OpenStudy (anonymous):

If x varies directly as y, and x=14 when y=2, find x when y=8. x=?

OpenStudy (imstuck):

y = kx is the formula. Replace y with 2 and x with 14 to find k.

OpenStudy (imstuck):

2 = 14k so k = 1/7

OpenStudy (imstuck):

Now use that k value in the same formula along with y = 8. 8 = 1/7x x = 56

OpenStudy (anonymous):

Thank you.

OpenStudy (imstuck):

You bet!